OpenStudy (anonymous):

4 log x-6 log (x+2) write the expression as a single logarithm. a.24 log x/x+2 b.log x^4 (x+2)^6 c.logx(x+2)^24 d.none

OpenStudy (anonymous):

\[\log\left(\frac{x^4}{(x-2)^6}\right)\]

OpenStudy (anonymous):

K so B.